Daily Plans



 

March 8, 2010

Today: Run Physics.
 

February 19, 2010

Today: One short access, then run Physics.
At 10 am we will make an access to check the vacuum around the AC dipole. In addition:
  1. LLRF will load new code to fix problem with transition timing slip.
  2. Stochastic cooling will make short access
  3. STAR magnet filter bags will be replaced
  4. STAR will make a 15 min. access.
  5. PHENIX will make a 15 min. access.
 

February 8, 2010

Run Physics, commission stochastic cooling.
  1. keep intensity >1.3e9 ions/bunch at injection
    But don't spend too much time tuning injectors to get higher intensity.
    We want to keep time between stores less than 1 hour.
  2. 111x111 bunches for all stores
  3. run with stochastic cooling; "coolers" may be around for development
    Blue longitudinal is now working. Use for physics.
  4. Rebucketing setup:
    • YS1 through YS3 set to 460 kV
    • BS1 through BS3 set to 800 kV
    • X2,X3,& X4 set to 800 kV
    • YS1, YS2,& YS3 keep raising by 100kV/store, to 800kV, if possible. Watch Yellow and Blue losses.
    • X1 off.
 

February 5, 2010

Today: Run physics, commission stochastic cooling
    From noon to 2pm will look at transition. (machine development).

This weekend: Run Physics.
  1. keep intensity >1.3e9 ions/bunch at injection
    But don't spend too much time tuning injectors to get higher intensity.
    We want to keep time between stores less than 1 hour.
  2. 111x111 bunches for all stores
  3. run with stochastic cooling; "coolers" may be around for development
  4. Rebucketing setup:
    • YS1 through YS3 set to 460 kV
    • BS1 through BS3 set to 800 kV
    • X2,X3,& X4 set to 560 kV
    • X2,X3,& X4 keep raising by 100kV/store, to 800kV, if possible. Watch Yellow and Blue losses.
    • X1 off.
 

February 4, 2010

Tonight/Overnight: Run physics.
  1. keep intensity >1.3e9 ions/bunch at injection
    But don't spend too much time tuning injectors to get higher intensity.
    We want to keep time between stores less than 1 hour.
  2. 111x111 bunches for all stores
  3. run with stochastic cooling; "coolers" may be around for development
  4. Rebucketing setup:
    • YS1 through YS3 set to 460 kV
    • BS1 through BS3 set to 800 kV
    • X2,X3,& X4 set to 460 kV
    • X1 off.
 

February 3, 2010

Tonight: APEX ends at 7pm, run physics.
  1. keep intensity >1.3e9 ions/bunch at injection
    But don't spend too much time tuning injectors to get higher intensity.
    We want to keep time between stores less than 1 hour.
  2. 111x111 bunches for all stores
  3. run with stochastic cooling; "coolers" may be around for development
  4. Rebucketing setup:
    • YS1 through YS3 set to 460 kV
    • BS1 through BS3 set to 800 kV
    • X2,X3,& X4 set to 460 kV
    • X1 off.
Tomorrow:
We will develop a new ramp with the STAR magnet polarity reversed. This will require doing a tune feedback ramp and a feed-forward ramp to compensate for the slight tune shift caused by the magnet.

Here is the plan for the day for tomorrow:
At 8:00 am we end the physics store, ramping down the STAR magnet.
  1. It will take the CAS watch a total of 3 hours to reverse the polarity and return the magnet for use. No access is required.
  2. During the 2 hours with the magnet off we will work on the following things (I assume 1/2 hour ramp up/down time):
    1. 8:30 am: Look at Blue injection efficiency, which has been poor
    2. 9:00 am: Do one 12 bunch ramp with tune feedback to commission new LLRF dsp code to keep beams cogged on the ramp.
    3. 9:30 am: The instrumentation group will look more at orbit feedback, to prepare for possibly using orbit feedback when the STAR magnet is back on and for the full ramp development.
  3. 10:30 am: magnet should be available again and can be ramped up.
  4. 11:00 am: 12 bunch ramp with STAR magnet on, with tune feedback and possibly orbit feedback
  5. Next ramp is a feed-forward ramp, if all went well. It is possible we will need two tune feedback ramps.
  6. The plan is to get back to physics by 6 pm. The intervening time will be spent getting back to full intensity, as best as possible, and looking at improving the transition crossing, if there is enough time.
As I implied in step 7, if the afternoon ramp development goes well we will take two hours for the RF group to look at cavity phasings. This would be to fix a problem we have with bunch lengths increasing at transition. For this work they will need to look at beam at injection, then a few 12 bunch ramps, and then look at phasings at store, including the rebucketing cavities. If this work isn’t done tomorrow afternoon then we will take a couple hours Friday morning for it.

Greg Marr will be the shift leader for tomorrows efforts.
 

February 1, 2010

Tonight: Run physics and stochastic cooling commissioning.
  1. keep intensity >1.3e9 ions/bunch at injection
    But don't spend too much time tuning injectors to get higher intensity.
    We want to keep time between stores less than 1 hour.
  2. 111x111 bunches for all stores
  3. run with stochastic cooling, both blue and yellow,if available
  4. Rebucketing setup: All cavities, except X1, at 460 kV, no ramping.Keep X1 off.
 

January 29, 2010

Today: Run physics and stochastic cooling commissioning.
  1. LLRF group will take 1/2 hour to work on injection timing when the morning store is finished (~10 am).
  2. Please try to get intensity back up >1.2e9 ions/bunch. The stores since maintenance ended have all been low.

  1. First store tonight (11578) is with all 200 MHz cavities at 460 kV each.
  2. For the next store (11579) turn off X1 and run the other 200 MHz cavities at 460 kV each.
  3. For the store after that (11580) turn off X1 and X2 and run the other 200 MHz cavities at 460 kV each.
  4. First the rest of the weekend use all 200 MHz cavities at 460 kV each.
This weekend: Run physics.
  1. keep intensity >1.2e9 ions/bunch
    But don't spend too much time tuning injectors to get higher intensity.
    We want to keep time between stores less than 1 hour.
  2. 111x111 bunches for all stores
  3. run with stochastic cooling, both blue and yellow,if available
  4. Rebucketing setup: All cavities at 460 kV, no ramping.
 

January 28, 2010

Today: Run Physics and Stochastic cooling commissioning.
 

January 26, 2010

Today: Run Physics and Stochastic cooling commissioning.

Tonight/Overnight: APEX

Tomorrow: Maintenance.Schedule

 

January 25, 2010

Today: run physics.
  1. Use yellow transverse cooling, but reduce amount of cooling after one hour into the store (after physics start) by increasing the "fear" factor from 3 to 10.
  2. For the next store (around noon today), we will try bringing on one of the common cavities, to see if we can imrove rebucketing.
Tonight/Overnight: run physics.
  1. Use yellow transverse cooling, reducing the amount of cooling after 45 minutes into the store by increasing the "fear" factor from 3 to 10.
    • At a factor of 10 there may still be too much cooling later in in the store. There are two things to try in this case.
    • First: turn off the cooling entirely when the yellow emittances get below 8 to 10 pi.
    • Second: on the next store, start out with a higher factor than 3 (perhaps 4, next store could try 5).
  2. Rebucketing strategy:
    • For the next store (after the current one ending at 17:47), raise the voltages on the 200 MHz cavities (3 main cavities/ring plus the X1 common cavity) to 700 kV/cavity, or a total of 2.8 MV.
    • If rebucketing is successful at 2.8 MV (look at losses and vacuum), try turning on another 200 MHz cavity in the store following the next one.
      For each successive store, add another common cavity until all four are in use.
    • If rebucketing is not successful at 2.8 MV, we should try to debug why. So the thing to do is rebucket with 5 cavities (so use two common cavities) keeping the total at 2.1 MV.
      For the next store use 6 cavities, again keeping at 2.1 MV.
      NOTE: WE ONLY DO THIS IF THE FIRST ATTEMPT AT 2.8MV FAILS.
    • The goal is to get all 200 MHz cavities on and running at store.
 

January 22, 2010

Today: get a physics store up on Au104 with rebucketing.
Tonight/Weekend: Run physics.
  1. Run all weekend with rebucketing but without the common 200 MHz cavities.
  2. For the next store raise the intensity/bunch up to 1.2e9
  3. For the store after that step up the number of bunches (93 or so).
  4. Next go to 103 bunches.
  5. If you can, run the rest of the weekend with 111 bunches and try to push up the intensity/bunch
  6. The goal is to integrate as much luminosity as possible this weekend, so we want to keep to the 4 hour stores and keep the time between stores minimal.
  7. Hopefully, Stochastic coolers can give us yellow transverse for the weekend.
 

January 21, 2010

This morning: run physics and prepare for machine development.
Today/Tonight: Machine Development: Au104
    Develop new ramp with 0.7m beta-star.
  1. use 6x6 ramp with tune feedback on Au104
    • include new horiz. orbit offsets for bpm's.
    • include yellow dipole feedback
  2. If it makes it to store, measure chromaticity and check radial aperture.
  3. We may need a 2nd tune feedback ramp.
  4. 2nd or 3rd ramp is the feedforward ramp.
  5. For each of these ramps the orbit needs to be corrected. Also for each that makes it to store, do step 2.
  6. Next do a wiggle ramp to measure chromaticity on the new ramp - and correct it.
  7. If all has gone well, go to 12x12 on the next ramp, look at efficiencies, and try rebucketing at store.
  8. Assuming success, increase number of bunches, correct orbits, and work towards getting a physics store with rebucketing.
    • progress to 28x28 bunches, then 54x54 - if it looks good, try 111x111
Overnight: Run physics, if ready, otherwise continue to develop Au104.
 

January 20, 2010

Today: APEX Schedule
Tonight/overnight: Run Physics.
 

January 15, 2010

Physics, with one rebucketing test ramp:

  • Ramp 23 bunches as for a physics store with rebucketing enabled once current store is dumped (~11:00 am)
  • Collimation will be in on the ramp as for physics (pending a check by KAD)
  • No common cavity will be used.
  • IR optimization may be necessary (post rebucket).
  • Assess possibility of success with 111 bunches based on instantaneous losses at rebucketing + lifetime. Decide after the test ramp and try ONE 111x111 store rebucketed. If this failes, do not try again but return to physics w/o rebucketing. If it succeeds, keep running physics with rebucketing.


  • January 14, 2010

    - Today: Run physics and stochastic cooling commissioning.
    - Tonight/Overnight: run physics. Stay with 4 hour stores.
           Do not spend more than 1/2 hour after a dump working on Injectors
           intensity to meet ions/bunch goals. Take the best possible and try
           to keep time between stores minimized.
           Check that injectors intensity is good at least 1/2 hour before dump.
     
           There will be one test performed at the end of the current
           store (as of 5pm). If this test is successful the next store
           will attempt rebucketing without the common cavities.
           If this works, we will run overnight in that mode.

           The test involves shifting the target frequency at store by +13 Hz
           and adjusting the Brho frequency offset by the same amount. This should
           center the yellow beam in the momentum aperture, based on previous
           measurements.
           Next, the radius will be shifted by small amounts to probe how much
           momentum aperture there is with the new frequency.
           If there is +/- 0.5 mm space without beam loss then that tests how
           much room the rebucketing gymnastic requires.
           If there is enough room, a rebucketing attempt can be made (still end of store).
           Nevertheless, on the next store, if there is +/- 0.5 mm, then rebucketing
           will be attempted, without the common cavities.


    January 13, 2010

    - Today: Maintenance until 6 pm. Schedule
    - Tonight/Overnight: run physics. Stay with 4 hour stores.
           Do not spend more than 1/2 hour after a dump working on Injectors
           intensity to meet ions/bunch goals. Take the best possible and try
           to keep time between stores minimized.
           Check that injectors intensity is good at least 1/2 hour before dump.
     

    January 12, 2010

    - Today: run physics, stochastic cooling commissioning.
    - Tonight until 8 pm: run physics
           Do not spend more than 1/2 hour after a dump working on Injectors
           intensity to meet ions/bunch goals. Take the best possible and try
           to keep time between stores minimized.
           Check that injectors intensity is good at least 1/2 hour before dump.
    - Overnight: APEX
     

    January 11, 2010

    - Today: run physics, stochastic cooling commissioning.
    - Tonight & Overnight: run physics
           Next store, push for 1.2e9 ions/bunch average
           For each new store try to increase ions/bunch average by 0.1e9

           Do not spend more than 1/2 hour after a dump working on Injectors
           intensity to meet ions/bunch goals. Take the best possible and try
           to keep time between stores minimized.
           Check that injectors intensity is good at least 1/2 hour before dump.
     

    January 8-10, 2010

    - For the weekend the plan is to run stores for physics.
       Saturday morning there will be a 2-4 hour access for
       both experiments and stochastic cooling folks. The
       access will start at 8 am, but may slip a couple hours
       depending on when the physics store is set to end.

    - Today: working on resolving why we can't turn on rebucketing without losing beam.
       1. Main issue seems to be negative chromaticities (and large) at store.
       2. So first measure yellow sextupole polarities (blue were already checked and okay).
       3. Next, make sure working points and chromaticity settings are correct.
           Blue Tunes: (0.237, 0.229)
           Yellow Tunes: (0.225, 0.235)
           All Chroms: small and > 0 (small, meaning what gives best lifetime)
       4. With small Chromaticity test radial range - how much loss is observed for +/-0.5 mm swing.
           At full voltage this is the equivelent range for the increase in dp/p
       5. With corrected store (low lifetimes, small chrom., etc.), test rebucketing at low intensity
           0.5e9 ions/bunch with 12 to 54 bunches, but not a full fill
       6. If all goes well, put up a physics store with rebucketing.
       7. If not, put up a physics store without rebucketing
     
    If rebucketing is successful, run the weekend physics stores with it, and work on losses and increasing luminosity.
    Each new store should add more bunches, with intensity up to 1.1e9 ions/bunch (average for all).

    If rebucketing is not successful, run the weekend without it, raising intensity for each new store.

    -Tonight; work until midnight on getting rebucketing to work, but have a physics store up by midnight.
     

    January 7, 2010

    - Today: morning access & afternoon beam development
       1. 2 hour access starting around 8:45 am
       2. beam development - automate start of store
          take two 6 bunch stores to redo end of ramp (move to Au103, with added stone)
          include IR corrections and proper chromaticities
          Do feedback ramp for first store, measure store chromaticity
          Second ramp with replay, adjust lifetime with chromaticity and tune scans
          Do a third ramp, putting up as physics store
       3. fix 5% loss in Blue snapback
       4. injection losses are high (>10%)
       5. drift compensation
       6. move working points away from 9th order resonance
       7. In physics ramp, tune lifetime with chromaticity and small working point changes.
    - Tonight: finish work from the day = need to have Au103 with correct store tunes and chromaticities.
    - Overnight: Run Physics on Au103, continue push for higher luminosity.
     

    January 6, 2010

    - Today: pay my property taxes.
       1. Run stores for Stochastic cooling and physics.
       2. first APEX tonight, starting either 4 or 6 pm.
     

    January 5, 2010

    The last few stores (11365, 11368, 11369, & 11370) showed very bad lifetimes
    in Yellow, so we chose this morning to use the day (actually 1/2 the day,
    which turned into the full day ... and some) to measure the chromaticities
    using 6 bunch ramps. The RF group also needed time to fix rebucketing, the
    injection cogging, and to get the Brho loop to work as is needed to set the
    right frequency at store.

    - Today: 6x6 bunch ramps to measure chromaticity, nonlinear chromaticity,
       fix rebucketing, injection cogging, and Brho loops.
    - Tonight:
       1. RF group will put in df for rebucketing frequency offset
       2. RF group will work at injection to fix injection cogging
       3. Parasitically; BPM timing checks, injection kickers fine timing
       4. 6x6 bunch ramp with Yellow Vert. Chrom. fixed at the store stone (+10 to 12 setpoint units).
       5. RF group will test if Brho loops work, with fix put in.
       6. At end of this ramp, look at phase errors in rebucketing
       7. Put up a 68x68 bunch store with 1.1e9 ions/bunch
       8. Tune yellow chromaticity to improve lifetime
    - Overnight:
    Store lengths should now go to 4 hours, with max. length at 6 hours.
    After 3 hours check with experiments to agree on end of store time.
       1. Next store, 68x68 with 1.2e9
       2. Next store, 78x78 with 1.2e9
       3. Next store, 88x88 with 1.2e9
       4. Umm? You get the idea, right?
     

    January 4, 2010

    - Today:
       1. current store stays until 1 pm
       2. Access for STAR and Stochastic cooling at 1 pm (1/2 hr)
       3. Put up another store, increasing bunch intensity.
       4. Stochastic cooling commissioning this afternoon (behind store).
    - Overnight:
    - Continue running stores, increasing intensity with each new store.
        (Use guidance per Jan. 1 entry below, for length of stores).
     

    January 3, 2010

    - Continue running stores, increasing intensity with each new store.
        (Use guidance per Jan. 1 entry below, for length of stores).
       1. For the next store go to 0.9e9 ions/bunch, keeping to 56x56 bunches.
       2. For each new store, increase bunch intensity by 0.1e9, until you either get up to 1.2e9, or hit a limitation (instabilities after transition, for example).
       3. After hitting the bunch intensity limit, cut back by 0.1e9 and then increase bunch number by 10, for each new store, up to 103x103.
       4. Work on losses, collimation, and use gap cleaner from beginning of store
       5. Consult with LLRF group and see if we can fill using 1st and 4th bunches in AGS (so we speed up the fill time by x2).
       6. Move yellow away from 9th order resonances (and vert. tune of 0.2) by raising both tunes = tune on yellow lifetime at store.
       7. For each new ramp, adjust chroms during beta-squeeze (follow what Joanne did in the day shift today) to reduce losses at end of store.
     

    January 2, 2010

    - continue running stores and increasing intensity, when possible
    (see Jan. 1 entry below).
    - Area's that seem to need work.
       1. Put collimator's to STBY on ramp (done, in tape, by Greg M.)
       2. understand lurches in corrector settings during beta-squeeze
       3. reduce losses in STAR triplets during beta-squeeze
       4. work on Brho loop and moving to correct frequency at store
       5. Fix Blue injection efficiency ( >5% worse than Yellow)
       6. Yellow transition loss - dipole trim correction enough?
       7. Get Yellow orbit feedback during transition working (Carl S.)
     

    January 1, 2010

    - Over the weekend: run stores for physics.
       1. Run stores for 6 hours, continuously over the weekend
           Negotiate with experiments for when to end stores,
           but don't keep any stores for more than 8 hrs
           Check with STAR and PHENIX 4 hours into store, to decide when to end it.
           Endeavor to keep time between stores minimized.
       2. For each new store, keep bunch intensity at ~1.0e9
       3. increase bunch number by 10, for each new store, to get to 103.
       4. Work on losses, collimation, and use gap cleaner from beginning of store
       5. Keep 1/2 hr orbit correction on during stores.
       6. Keep IPM on during ramp and stores
       7. Once doing continuous stores, start automatic orbit correction (requires 24 hours stores)
       8. Still need to finalize beginning of store progression, with extra stone (Greg M.)
     

    December 31, 2009

    - Today: commissioning LLRF for storage cavities.
       1. in parallel, stochastic cooling can use beam at store.
    - Tonight: when rebucketing is setup and RF group is finished
       1. setup a store with 23x23 bunches at 0.7-0.8e9 ions/bunch for automatic collimation development.
       2. next store: setup with 56x56 bunches at 0.7-0.8e9 ions/bunch.
    - Tonight & Over weekend: begin to run physics
       1. Run stores for 4 to 5 hours continuously over the weekend
       2. For each new store, increase bunch intensity by 0.1e9, until you either get up to 1.2e9, or hit a limitation (instabilities after transition, for example).
       3. After hitting the bunch intensity limit, increase bunch number by 10, for each new store, to 103.
       4. Work on losses, collimation, and use gap cleaner from beginning of store
       5. Keep 1/2 hr orbit correction on during stores.
       6. Keep IPM on during ramp and stores
       7. Once doing continuous stores, start automatic orbit correction (requires 24 hours stores)
       8. Still need to finalize beginning of store progression, with extra stone (Greg M.)
     

    December 30, 2009

    - Today/Tonight: Setup and commission rebucketing with storage cavities.
       1. use 6x6 ramp with 0.8e9 ions/bunch
       2. yellow replay function needs smoothing
       3. Effort is to end in time to have store up at midnight for exp. setup.
    - Overnight: Put up a physics store for experiment setup. 56x56 with ~0.8e9 ions/bunch.
    - Tomorrow morning: if LLRF needs more time for storage cavities, continue commissioning.
       1. PS group needs 1/2 hr to work on Yellow quad mains PS - no beam.
       2. if LLRF is finished, continue Stochastic cooling effort.
     

    December 29, 2009

    - Tonight: work on store development, increase bunch intensity
       1. Work toward >1.0e9 ions/bunch and Nb>56
       2. Look at bpm timing (Todd)
       3. Work on collimation and reduce background rates
       4. check/fix coupling correction at store
       5. At end of store: do vernier scan (crossing angle and vernier scans)
       6. At end of store: chromaticity measurement (Yun)
    - Overnight: Put up a physics store for experiment setup.
       -- best intensity/bunch and bunch number based on evening work
    - Tomorrow morning: 6 bunch store(s) for development
       1. Stochastic cooling commissioning
       2. Optics Measurement (Vadim)
       3. Work on HL drive for storage cavities (phase adjustments)
       4. Chromaticity on ramp
    - Tomorrow day: Back to physics store for rebucketing and stochastic cooling development
     

    December 28, 2009

    - Today: Keep a store for exp. setup and stochastic cooling development.
    - Tonight: LLRF work.
       1. turn on Brho loops
       2. match injection to store bunch pattern (fill injection into collision pattern)
       3. RHIC filling problem - only 1 bunch/cycle when > 23 bunches?
       4. why the gap in Blue bunch pattern?
    - Tonight: beam development work
       1. push bunch intensity limit (>1.1e9?)
       2. can we get to 86 bunches
       3. measure longitudinal emittance at store
       4. chromaticity on blue ramp
       5. fix coupling at store
    - Overnight: put in store for experiment setup
    - Tomorrow: continue running stores
       1. Stochastic cooling development.
       2. Set store sequence at beginning of store (Greg M.)
     

    December 26, 2009

    - Today/Tonight: Get Ring to Ring Synchro and store cogging working.
       1. Also get the Yellow Landau cavity on (Blue should already be on).
       2. Begin working on setting up collisions, with proper timing in place.
       3. Both Yellow and Blue should be running on replay.
       4. Yellow will need to be retuned now it is slave to Blue.
    - Overnight: Can we put in a first store for the experiments?
       1. Setup store sequences
       2. Look at gap cleaning
       3. collimation?
    - Tomorrow: Increase bunch intensity and bunch number
       1. bunch intensity limit? >1.1e9?
       2. can we get to 64 bunches tomorrow?
     

    December 24, 2009

    - Today/Tonight: With both beams getting to store, begin working on:
       1. opening RF loops and running on fixed frequency at store.
       2. tune store beam lifetimes - orbits, chromaticities
       3. put beams at store into collisions - manually
       4. get both rings running on replay with separation bumps on.
       5. measure chromaticities on ramps
       6. measure beam optics at store (ac-dipole measurement)
       7. measure dispersion, chromaticity, IPM at store.
    - for the next couple of days
       1. improve ramp/store efficiencies
       2. raise bunch intensities up to > 1.1e9
       3. LLRF commissioning = ring to ring synchro, Yellow Landau cavities, store cogging.
       4. begin working on increasing number of bunches.
     

    December 23, 2009

    - Today: Focus on transition crossing in both Blue and Yellow.
    - Tonight: Work on Yellow ramp efficiency
       1. Transition crossing - correct orbit
       2. LLRF will work on Ring to Ring synchro
       3. correct orbit through beta-squeeze into store
       4. turn on separation bumps (if not on)
       5. once good efficiencies, put Yellow on replay.
       6. Measure chromaticities on the ramp.
       7. Blue: on replay, measure chromaticities.
    - Overnight: continue working on Blue/Yellow efficiencies.
       1. work on establishing collisions (manual cogging only for now).
       2. if losses on ramp into store are low, increase bunch intensities
       3. Look at collisions on ZDC's = check that HV PS's are on.
     

    December 22, 2009

    - Today: Focus on bringing beam to store in both Blue and Yellow
       1. Tune beam past the loss at t220-t240, in Blue.
       2. Work on transition losses in both rings.
    - Tonight: Continue work on ramps, bring beam to store.
       1. Turn on separation bumps in both rings. Ramp down from 5 mm to 3 mm during beta-squeeze.
       2. Get Blue running on replay, if possible.
       3. If Yellow is in good shape, get running on replay.
       4. Measure chroms on ramp - tune to reduce coherence.
    - Overnight: Continue ramp setup, tune efficiencies on ramp an store.

    Note: Ring to Ring synchro will not be ready to use until tomorrow. Don't expect collisions overnight.


    December 21, 2009

    - Today: Ramp/Store development in both Blue and Yellow.
       1. A problem has developed in Blue, beam is lost at t220. Diagnose this problem.
       2. LLRF commissioning in Yellow. Get Loops working.
           note: RhicInjection needs phase detector data.
       3. Yellow ramp needs work - Quads max out leaving no room for Tune feedback to correct. Steve T. is working on it.
       4. There is a laundry list of items developing:
           Blue - 20 bunch gap problem
           Landau Cavities
           Quad mode damper at transition.
           Cogging/rebucketing for collisions.
           Ramp/store measurements: chromaticity and dispersion.
       5. When beam is at store, we need to give Stochastic cooling some time to look at things.
     

    December 20, 2009

    - Tonight: As soon as beam is available
       1. Re-establish Blue with 6 bunches - ensure beam still gets to store.
       2. Yellow is now available to ramp on Au102. Begin working on acceleration in Yellow.
           Use tune/coupling feedback and use orbit feedback, if possible, to get through transition.
           As with Blue, you will very likely have to manually correct the orbit during the beta-squeeze.
           Re-connect the wfgman's = Al left instructions at 23:13 in travis's (rhic-au_2010-Sat_Dec_19_2009_20:49:44_PM) elog.
     

    December 18, 2009

    - Today/Tonight: work on Blue ramp/store development.
       1. Work on Blue with Tune/coupling feedback, but with no orbit feedback.
          a. The first game is to understand the 1 mm offset during the ramp (RF group).
          b. Note, the RF group will also correct the 5 nsec offset, fixing the instrumentation timing offset.
          c. Turn on Blue separation bumps.
          d. Develop Tune/coupling replay with separation bumps on.
          e. Check orbit and make any corrections manually.
          f. Measure ramp chromaticities.
       2. Begin working on Blue store corrections and measurements.
           Store dispersion and chromaticities.
       3. LLRF will continue to look at cogging and other issues.
       4. Ramp/Store development: push intensity/bunch (>1e9), then increase #bunches and tune losses.
       5. Yellow MMPS is not ready, PS group will continue to work on it today.

    - Overnight (eve & owl): If/when Yellow is available begin Yellow ramp/store development.
       1. Work on Yellow with Tune/coupling feedback and with orbit feedback.
       2. Correct orbit on the ramp and get beam to store.
       3. When Yellow gets in good shape, bring on separation bumps and develop replay ramp.

    - Once we have both Blue and Yellow with beam at store we can begin working on colliding beams.
    - The weekend goal is to get to collisions.  

    December 17, 2009

    - Today and this evening: work on Blue ramp/store development.
       The focus is on the magnet transfer functions and getting a model that doesn't get driven into the integer tune.
       So a new ramp was developed using the old transfer functions (they were changed for this run - and appear to be a problem) which will be worked on this evening.
       The new ramp is called Au102 (old one was Au101). It has all the same features, but referenced to the old magnet transfer functions.
       
       In addition, continue setup of Yellow MMPS.
       Yellow was off all day while the PS group worked on a QPA problem (either in Q6 or Q7 ps).
       As of the end of the day, they still have not fixed the problem.
       So Yellow is still not ready for ramp/store development with beam.
       
    - Overnight (eve & owl): Continue working on Ramp development.
       1. Yellow MMPS is not ready, continue work on Blue ramp development.
       2. Ensure experimenter magnets are on.
       3. Work on Blue with Tune/coupling feedback, but with no orbit feedback.
           Turn on Blue separation bumps.
           Develop Tune/coupling replay with separation bumps on.
           Check orbit and make any corrections manually.
           Measure ramp chromaticities.
       4. Begin working on Blue store corrections and measurements.
           Store dispersion and chromaticities.
       5. LLRF will continue to look at cogging and other issues.
       6. If cogging is working, push intensity/bunch (>1e9) as high as possible and fill at injection.
           This is an injection only test, to see how the vacuum responds.
       7. Ramp/Store development: push intensity/bunch (>1e9), then increase #bunches and tune losses.
       8. Orbit feedback: try fixed orbit seperation bumps with orbit feedback, if ready.

    - Tomorrow: Continue LLRF commissioning, both Blue and Yellow ramp development.
     

    December 16, 2009

    - Today and this evening: repair the AGS Motor Generator oil leak.
       In addition, continue setup of Yellow MMPS

    - Overnight (late eve & owl): Return to working on Ramp development.
       1. Yellow MMPS is not ready, cont. work on Blue ramp development.
       2. Re-establish Blue and Yellow injection and RF capture.
           Ensure experimenter magnets are on.
       3. Continue working on Blue orbit corrections.
       4. Begin working on Blue store corrections and measurements.
           Store dispersion and chromaticities.
       5. LLRF will look at cogging and other issues.
       6. If cogging is working, push intensity/bunch (>1e9) as high as possible and fill at injection.
           This is an injection only test, to see how the vacuum responds.
       7. Ramp/Store development: push intensity/bunch (>1e9), then increase #bunches and tune losses.
       8. Orbit feedback: try fixed orbit seperation bumps with orbit feedback, if ready.

    - Tomorrow: Continue LLRF commissioning, both Blue and Yellow ramp development.
     

    December 15, 2009

    - This evening/overnight: Continue Blue Ramp development.
       1. The goal is to get to store and then work on ramp efficiency.

    - Tomorrow: Off all day to work on AGS Motor Generator, which has an oil leak.
       1. No beam development from 4 am to midnight.
       2. Limited access to RHIC tunnels, since MMPS work continues for both rings.
       3. Exp. areas can be open.


    December 15, 2009

    - This evening/overnight: Continue Blue Ramp development.
       1. The goal is to get to store and then work on ramp efficiency.

    - Tomorrow: Off all day to work on AGS Motor Generator, which has an oil leak.
       1. No beam development from 4 am to midnight.
       2. Limited access to RHIC tunnels, since MMPS work continues for both rings.
       3. Exp. areas can be open.


    December 14, 2009

    - Today: Preparing Blue for ramp development.
       1. PS group should be done by 2:30 pm, and we can begin Ramp development.
       2. In parallel, PS group will work on preparing Yellow to run.
       3. Starting with the first ramp, we will now run with exp. magnets on.
       4. The main goal is to commission acceleration with new LLRF.
    - This evening: Continue Blue ramp development.
       1. LLRF commissioning acceleration, transition phase jump, to store.
       2. PS group will continue to work on preparing Yellow to run.
       3. BBQ and Tune/coupling feedback is on for every ramp.
         (in principle we can use replay - without interfering with LLRF)
         Remember we cycle off the Booster during ramps with tune feedback.
       4. Setup Gamma-tr jump.
       5. If/when possible, work with orbit feedback on the ramp.
    - Overnight: Continue Blue ramp development.
       1. Assuming acceleration is finished, begin working on ramp efficiency
          increase single bunch intensity (up to 1e9 ions/bunch).
          increase number of bunches, if possible (need injection cogging).
       2. Setup instrumentation = ipm's, can we get any emittance data?
       3. Chromaticity measurment, non-linear chrom.??


    December 13, 2009

    - This evening: LLRF is working on commissioning in Yellow.
       1. In parallel work on instrumentation (BPM's, WCM, BBQ, BTF, ...)
       2. Check and correct orbits.
       3. Measure tunes and set to correct values.
    - Overnight: Continue work in Yellow.
       1. Test tune feedback.
       2. Checkout orbit feedback in Yellow.
       3. Measure Chrom. and dispersion
       4. Check injection match and Y-arc tune.

    - Tomorrow day: begin working on Blue acceleration, to Store.
       1. At 7 am, PS group needs access to RHIC to fix bi8-qgt-ps.
       2. At 8 am Don Bruno will start preparing Blue to ramp at the normal rate.
       3. When the gamma quad is fixed, Don can try a ramp.
       4. By late morning we can begin working on acceleration in Blue.
       5. In parallel, Carl S. will be working on Yellow MMPS, preparing to ramp.


    December 12, 2009

    - Tonight: work on Yellow beam at injection. LLRF in Yellow (Capture).
       1. Re-establish injection and work on circulating beam.
       2. Probe apertures around 12 O'Clock and Yellow cold snakes.
          These were areas of losses earlier today.
          Note: we work without hysteresis loop in Yellow.
       3. Correct Yellow orbit = try Run9 correctors for Yellow.
       4. While work in Yellow, PS group will work on Blue. (Au100)
          If they need to ramp, yellow beam work is suspended.
       5. Contact RF group when there is circulating beam.
       6. Checkout yellow instrumentation, BBQ, BTF, etc.
       
    - Tomorrow: work with Blue ramp development - may have reduced acceleration rate.
       1. PS group may not get the Blue ramp to run at the full rate of Au100.
           They will leave us with the best they can develop.
       2. While beam development in Blue, PS group can work on Yellow Ramp.
           Carl needs most of the day and evening to study Yellow.
           Yellow MMPS has new regulator and PLL, as does Blue, but
           it also has the new auto-transformers.
           So Carl needs to spend a good amount of time on Yellow.
       3. For Blue, work on LLRF commissioning = acceleration.
           Work with whatever ramp rate PS group suggests.
           BBQ and tune feedback for every ramp.
           If LLRF agrees, or is done, can try replay. (okay, the experts decide what is best for keeping beam on the ramp.)
           Transition phase and gamma-tr jumps need to be setup.
           Remember, timing should be set for Au100 (adjusted for ramp rate?).
       4. As soon as Yellow is ready for ramps, begin Yellow acceleration work.
          If RF group is ready.
          Same drill: BBQ and tune feedback every ramp.
          Checkout Yellow instrumentation with beam.

    We will hold off reconnecting the WFG's for the two rings until we can settle the ramp rate issue. So for the weekend we remain decoupled and work on each ring individually. We don't need to put in the separation bumps yet, then.

    Sorry I can't set more definite time/shifts for the work, but it all depends on the hours the PS group works. So we are dependent on their work/sleep schedules at the moment.



    December 11, 2009

    - Today: work on Blue & Yellow MMPS
       1. PS group working on blue ramp.
       2. Yellow valve box work. Preparing Yellow for injection.

    - Tonight: work on Blue & Yellow MMPS
       1. For the Blue ramp, we are switching to Au100.

    The Au101 ramp contains a distortion of the Q6/Q7 quadrupoles in Blue to allow the gamma-tr jumps in Blue and Yellow to occur at different times. This distortion puts a "bump" in the currents around transition. The "bump" is causing difficulties in tuning up the MMPS's. The difficulties are significant enough that, to tune them out, it will require significant time (a couple of days) before we could actually run beam on a Blue ramp. The Au100 ramp does not contain this feature, but is otherwise identical to the Au101 ramp. So we have to give up on the separate gamma-tr jumps, for now, so we can get the beam developed and commission the LLRF.

       2. Yellow valve box work.

    The PS group is working hard to understand why one cable in the Yellow valve box is reading high resistance. The repair will go into the evening and depending on when they can fix it we may or may not get Yellow working tonight. We will have to stay tuned.

    - This weekend:

    The goal is to get beam to store in both Yellow and Blue by Monday morning. We will revisit the state of the two machines tomorrow at 3:30 and form a more detailed plan at that time.


    December 10, 2009

    - Tonight: no ramp in Blue yet
       1. PS group will work on Blue MMPS, possibly one ramp.
       2. Between any Blue ramps, work with beam at injection.
       3. Once Yellow is ready, work with beam at injection
        To get Yellow ready:
        == Need to finish polarity checks (5 pm)
        == There are tests to do at injection current
        == Need to demonstrate a hard shut-off.
       4. If working in Blue, focus on:
        == Measure Chromaticity
        == Measure Dispersion
        == Measure Sextupole Polarities
        == Test InjectionTuning application, with new LLRF
        == Test LLRF cogging
        == Test LLRF radial steering (needed for measurements)
        == Test LLRF Yellow synchro
        == Test filling the entire ring (with bunches from AGS), then dump
     This work will go on all night into the morning.
     Tomorrow we will evaluate when we can begin working on ramps in Blue.
     It is expected we can start tomorrow afternoon.

    - Friday:
       1. Assuming we can work on Blue ramp development, commission LLRF acceleration.
         Work on transition phase jump and gamma-tr jump
         Tune feedback for all stores = develop replay function.
       2. Begin working on Yellow MMPS preparation for ramp development.


    December 9, 2009

     - Tonight: two activities
        1. PS group is preparing Blue MMPS for ramps
        2. Between Ramps LLRF will continue setup at injection.
     - At midnight:
        1. PS work ends.
        2. Begin ramp development with beam, if MMPS is ready.
          or, work with Blue at injection
     - Tomorrow day:
        1. Finish all work in the tunnels.
        2. Prepare Yellow PS for injection
     - Tomorrow evening/owl:
        1. Work on yellow injection and capture
        2. When yellow is in good shape, return to blue ramp devel.
     - Friday day:
        1. Blue ramp development
        2. prepare yellow MMPS for ramp development.


    December 8, 2009

     - Today: bake-out work, p.s. check-out, and DX training
     - Evening: P.S. group works on preparing Blue for ramping.
        This work ends at ~8 pm.
     - After 8 pm: begin beam development again.
        1. Fix timing in AGS to RHIC transfer.
            AGS extraction kicker timing
            Instrumentation timing
            RHIC LLRF timing
        2. LLRF commissioning at injection field.
        3. Instrumentation checkout.
     - Over owl: continue LLRF commissioning
        1. LLRF
        2. if radial control is available
            (BBQ should be working)
            Dispersion measurement
            Chromaticity measurement
        3. Check sextupoles (Yun)
        4. Check coupling (Yun)
     
     - Beam work ends at 6 am, to get early start tomorrow.
     - Tomorrow, RHIC sweeps begin at 12 noon, so PS group can continue Blue checkout.


    December 7, 2009

     - Today: work on vacuum bake-out and p.s. check-out.
     - 4 pm: RHIC is locked up and begin DX training again.
     - Eve/owl: re-establish injection and circulating beam in Blue.
        scan aperture at 4 O'Clock, where there was a loss last night.
        commission LLRF (establish capture).
        correct injection orbit
        BBQ setup
        Injection Damper setup


    December 6, 2009

     - Still trying to keep to the Schedule.
     - Today and tonight work on DX training.
     - DX training is likely to go until ~2 am, Monday.
     - Over the owl shift work again on getting Blue injection setup.
       LLRF group is needed, since Blue Synchro is not setup.
       once beam is circulating, RF group can work on capture.
     - Injection work can go into the day shift, if needed.
       Vacuum group needs access at some point to check bakeout.
       PS group needs to work on Yellow sextupole PS's.


    December 5, 2009

     - Since no beam work was done last night, keep to Dec. 4 plan.
      Except: once beam is circulating, contact RF group and work on capture.
     - P.S. group is working on final tests for Blue.
     - Expect to be able to inject around 9 pm.


    December 4, 2009

     - First, power supply group needs to complete checkout of MMPS.
     - They should be done by 9 to 10 pm tonight. Then we can begin injection setup in Blue.
     - Start with the Run 7 corrector settings.
     - Tune the AGS field to get the rev. freq. right in RHIC (on WCM).
     - Setup the injection kickers and timing - get hundreds of turns (no RF).
     - Checkout 1st turn BPM's (call Todd, he will work from Home).
     - Over the owl, finished getting good circulating beam.
     - At 8 am, Saturday, hand over RHIC for vacuum and power supply work.


    December 3, 2009:

     - Blue is down to 4-5 K, starting cooldown of Yellow.


    December 2, 2009:

     - Cooldown continues on schedule. Blue < 10 K now.
     - Blue Main Power Supply being checked out. Not ready until Friday.


    December 1, 2009:

     - Start 4.5 K wave in Blue.
     - Prepared plan for first week of beam development; see Plan from Dec.1 to Dec.8
     - Beam has been extracted to the AtR beam dump as of Wed., Nov.25.


    November 19, 2009:

     - Finishing up Dry Run.
     - Today, LLRF begins switching over to new system.
     - Beam was injected and captured in AGS. Setup continues per Schedule.


    September, 2009:

     - working on run planning :)